OpenStudy (anonymous):

The top of a building in Ottawa is 104m above the ground. Suppose an object was thrown upward with an initial velocity of 19.6m/s from this height. The approximate height of the object above the ground, h metres, t seconds after being thrown, would be given by the function: h=-4.9t^2+19.6t+104 a) what was the maximum height of the object,above the ground? b) after how many seconds did the object reach its maximum height? c) From the time the object was initially thrown, how much did it takes to reach the ground, to the nearest second?

OpenStudy (anonymous):

Do you just need the answers? If you need the work shown, are you doing differentiation and derivatives to do this or other methods?

OpenStudy (anonymous):

i need the work shown and the ans please.

OpenStudy (anonymous):

and i think were usin the quad formula

OpenStudy (anonymous):

ohh, ok XD

OpenStudy (anonymous):

x-coordinate of the Vertex or the maximum height is found using this equation: -b/2a a=-4.9 b=19.6 -19.6/(2(-4.9)) =-19.6/-9.8 =2 x=2 for vertex Plug it back in to figure out the y-coordinate

OpenStudy (anonymous):

or t=2 in this case XD

OpenStudy (anonymous):

h=-4.9(2)^2+19.6(2)+104 =-19.6+39.2+104 =123.6 (2, 123.6)

OpenStudy (anonymous):

SO, the height is 123.6 and the time it took to get there is 2 seconds.

OpenStudy (anonymous):

that's the answer for the first two questions

OpenStudy (anonymous):

thanks :) - from charmi too

OpenStudy (anonymous):

welcome, now, for the third question right?

OpenStudy (anonymous):

but like we learned in school , that the "b" value is the maximum height thats the part we dont get :S

OpenStudy (anonymous):

just a sec

OpenStudy (anonymous):

I don't think that's right. The only way that would be true is if the coefficient of x^2 was 1/2.

OpenStudy (anonymous):

ohh, so like the question isnt right? for part (c) ?

OpenStudy (anonymous):

?? In question C, you just need to let h=0

OpenStudy (anonymous):

ohh ok thanks :)